OpenStudy (anonymous):

use the order of operations to evaluate the expression when x= 4, y = 2 , and z = 10 xy to the second power

OpenStudy (anonymous):

\[(xy)^2\] or \[xy^2\]?

OpenStudy (anonymous):

the second 1

OpenStudy (anonymous):

first one is \[(4\times 2)^2=^2=64\] second one is \[4\times 2^2=4\times 4=16\]

OpenStudy (anonymous):

thank u